please write answer clear,
Calculate the pH of 0.20 M Ba(OH)2 (aq) and 0.20 M NaOH (aq). should they be the same? why or why not?
Answer #1

They will not be same because concentration of OH- is different in both the cases.

For NaOH,

[OH-] = [NaOH]

= 0.20 M

For Ba(OH)2:

[OH-] = 2*[Ba(OH)2]

= 2*0.20 M

= 0.40 M

1) NaOH

use:

pOH = -log [OH-]

= -log (0.2)

= 0.699

use:

PH = 14 - pOH

= 14 - 0.699

= 13.301

Answer: 13.30

2) Ba(OH)2

use:

pOH = -log [OH-]

= -log (0.4)

= 0.3979

use:

PH = 14 - pOH

= 14 - 0.3979

= 13.6021

Answer: 13.60